OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 4x + 8 • x • 3 = 28x 1. 4x + 8 • x • 3 = 4x + 8 • 3 • x 1. ? 2. 4x + 8 • x • 3 = 4x + 24x 2. Multiplication 3. 4x + 8 • x • 3 = 28x 3. ? A. Associative Property of Addition, Addition B. Commutative Property of Addition, Multiplication C. Associative Property of Multiplication, Addition D. Commutative Property of Multiplication, Addition

OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 4x + 8 • x • 3 = 28x 1. 4x + 8 • x • 3 = 4x + 8 • 3 • x 1. ? 2. 4x + 8 • x • 3 = 4x + 24x 2. Multiplication 3. 4x + 8 • x • 3 = 28x 3. ?